### v1

@@ Abstract
Numerical simulations of detonation-containing flows have emerged as crucial tools for designing next-generation power and propulsion devices. As these tools mature, it is important for the combustion community to properly understand and isolate grid resolution effects when simulating detonations. To this end, this work provides a comprehensive analysis of the numerical convergence of unsteady detonation simulations, with focus on isolating the impacts of chemical timescale modifications on convergence characteristics in the context of operator splitting. With the aid of an adaptive mesh refinement based flow solver, the convergence analysis is conducted using two kinetics configurations: (1) a simplified three-step model mechanism, in which chemical timescales in the detonation are modified by adjusting activation energies, and (2) a detailed hydrogen mechanism, in which chemical timescales are adjusted through ambient pressure modifications. The convergence of unsteady self-sustained detonations in one-dimensional channels is then analyzed with reference to steady-state theoretical baseline solutions using these mechanisms. The goal of the analysis is to provide a detailed comparison of the effects of grid resolution on both macroscopic (peak pressures and detonation wave speeds) and microscopic (detonation wave structure) quantities of interest, drawing connections between the deviations from steady-state baselines and minimum chemical timescales. This work uncovers resolution-dependent unsteady detonation regimes, and highlights the important role played by not only the chemical timescales, but also the ratio between chemical and induction timescales in the detonation wave structure on simulation convergence properties.
@@ 1 Introduction
Numerical simulation of detonation waves in both canonical [1, 2, 3] and complex configurations [4, 5] is becoming increasingly important due to emerging interest in detonation-based propulsion devices [6, 7]. Such simulations are also crucial in other related applications, including condensed-phase explosives [8, 9] and hazard prediction [10]. Detonation dynamics in these applications are complicated by counter propagating waves, heterogeneous mixtures [11], and surface stabilization effects [12], leading to complex detonation structures influenced by fine-scale flow-chemistry interactions that are difficult to probe experimentally due to high pressures and speeds of the system.
Due to this inherent complexity, significant effort has been devoted to developing robust and predictive numerical methods for solving the compressible reacting Navier-Stokes equations, which govern detonation-containing flows. Recent approaches have utilized numerics rooted in finite volume methods for full-geometry [13] and canonical adaptive mesh refinement (AMR)-based simulations [14]. Additionally, higher-order extensions such as the piecewise parabolic method coupled with conservative shock tracking strategies [15], as well as discontinuous Galerkin methods [16], have been effectively used to study fundamental detonation dynamics. Alternative strategies like the ghost fluid method can also be used [17], which resolves shock-chemistry interactions through level set equations.
@@
Regardless of the approach used, proper understanding of grid convergence properties in PDE-based simulation of detonating flows is of utmost importance. Analysis of detonation convergence properties to this end has primary focused on identifying onsets of instability, with observations showing how acoustic waves generated in the induction zone at high resolutions – which would otherwise have been implicitly filtered at low resolutions – lead to detonation quenching behavior [19, 20]. Additionally, convergence analysis performed in Ref. [21] for ethylene detonations reports small variation in mean wave speeds, as well as the emergence of high-frequency and low-amplitude propagation modes in highly resolved cases. Similar behavior was observed in the context of n-Heptane detonations [22], where an initial convergence analysis step was carried out to isolate resolution effects on pulsating frequencies.
@@
As such, the goal of this work is to investigate the impact of chemical timescale modifications on the numerical convergence of unsteady detonations using simulations based on operator splitting. The objective is both to conduct a comprehensive detonation convergence analysis, and to isolate the chemical timescale as a tunable parameter from two different angles: (1) through direct modification of reaction mechanism parameters (i.e., activation energies), and (2) through modification of the simulation operating conditions (ambient pressures). The former leverages a simplified model three-step mechanism [24] and the latter leverages a detailed hydrogen mechanism [25]. More specifically, simulation convergence trends in terms of quantities of interest (QoIs) relevant to detonation engine design, such as wave speeds, peak pressures, and detonation wave structures, are analyzed using highly resolved simulations of unsteady detonations. Where applicable, comparisons are made to steady-state Zeldovich-von Neumann-Döring (ZND) baseline solutions, such that effects of chemical timescale modifications on convergence trends can be monitored through deviations of QoIs from ZND solutions in highly resolved regimes.
@@ 2 Numerical Approach
The compressible reacting Navier-Stokes equations are discretized in time using a globally second-order Strang splitting approach [26], which decouples the effects of chemical reactions from those of fluxes. The strategy used here is the routinely used “Reaction-Advection/Diffusion-Reaction” order of operator invocations. As such, there are two chemical time integration steps in a single global time step advance.
The equations are solved using an in-house solver based on the AMReX framework [27]. The solver is a block-structured adaptive mesh refinement (AMR) extension of the extensively verified UMReactingFlow [13]. On each refinement level, chemical time integration is handled using an adaptive explicit method, and the advection-diffusion advance is treated using a stability-preserving second-order Runge-Kutta method combined with a slope-limited approximate Riemann solver. This work specifically leverages a van Leer limiter with the Harten–Lax–van Leer contact (HLLC) approximate Riemann solver [28]. The reader is directed to Ref. [13] for additional detail on the numerics of each component.
Grid refinement is handled by AMReX [27], which provides a backend for the grid communication steps used in the block-structured formulation. The AMR tagging criterion is based on pressure and density gradients with additional buffer cell padding, and was determined to properly isolate the detonation wave structures for subsequent mesh refinement. Grid refinement factors were set to 2, and all convergence resolutions reported in Sec. 3 are extracted from the finest (highest) AMR level, with the highest level count being 14. Separate verification studies were conducted to ensure that AMR-induced interpolation errors did not pollute detonation profiles for the range of levels considered. Examples of unsteady detonation profiles showcasing AMR level coverage in a 12-level case for both simplified (Sec. 2.1) and detailed (Sec. 2.2) chemistry configurations is shown in Fig. 1. The figure shows that the refinement criteria used in the AMR simulations successfully resolve the detonation profile at the desired level specification.
@@ 2.1 Simplified Chemistry
There are four key parameters: the initiation and branching activation energies ($E_{I}$ and $E_{B}$), and the corresponding crossover temperatures ($T_{I}$ and $T_{B}$). This work leverages the activation energies to control the sensitivity of the detonating flow field to chemical reactions, providing a direct pathway to parameterize chemical timescales within the detonation wave structure, as showcased in Sec. 3.1. In particular, three different activation energy settings (Cases 1-3 in Table 1) were prescribed to explore a wide range of chemical timescale influence in the simplified configurations. Besides the activation energies, all other parameters match those used for the lowest $T_{B}$ case in Ref. [29]. To admit a theoretically valid detonation solution with this mechanism, the gas is assumed to be calorically perfect with constant specific heat ratio $\gamma$, leading to a simplified definition of chemical energy. This differs from the detailed case, where variable $\gamma$ is used. Additionally, viscous terms in the governing equations are neglected when using the simplified model mechanism. All length and time values in the simplified configurations are reported in non-dimensional quantities, where non-dimensionalization is carried out such that the termination rate $k_{C}$ is unity. The reader is directed to Ref. [29] for additional details.
@@ 2.2 Detailed Chemistry
Detailed chemical kinetics are modeled using the 14 species, 42 reaction H2-air mechanism of Shepherd et al. [25], which is based on the mechanism of Mével et al. [30]. The one-dimensional domain was 0.8 m long with a uniform base grid of 1600 cells, yielding a base resolution of 500 $\mu$m. The minimum grid size in the finest simulation is 30.5 nm. All cases used the reactant molar composition 2H2:O2:7Ar at an ambient temperature of 300 K. Significant argon dilution was added to ensure stability of the detonation waves [31]. In order to explore different chemical timescale regimes, ambient pressures of 0.2, 0.5, and 1 atm were investigated (as described in Table 1). For each case, the ZND solution was computed using the Shock and Detonation Toolbox [32]. To initialize unsteady simulations, this solution was then mapped onto the computational grid at the initial time, with the wave front positioned 0.2 m from the left boundary (i.e., the driver gas was the ZND wave in the detailed cases, as these were found to more quickly lead to self-sustained detonations).
@@ 3 Results
$\tau_{\text{chem}}/\tau_{\text{ind}}$ 1.218 ×$\times 10^{-1}$ 1.920 ×$\times 10^{-3}$ 5.448 ×$\times 10^{-5}$ 1.503 ×$\times 10^{-3}$ 1.631 ×$\times 10^{-3}$ 1.738 ×$\times 10^{-3}$
Table 1: Detonation case configurations. First row: activation energies in simplified mechanism normalized by squared speed-of-sound. Second row: ambient pressures. Third row: minimum chemical timescale observed in steady-state solution. Fourth row: Ratio of minimum chemical timescale and induction timescale in steady-state solution. Induction timescale defined as $\tau_{\text{ind}}=L_{\text{ind}}/w_{\text{CJ}}$, where $L_{\text{ind}}$ is induction length and $w_{\text{CJ}}$ is the CJ speed. ZND profiles shown in Fig. 2.
@@ 3.3 Convergence of Wave Structure
This finding illuminates how deviation from the steady-state solution in the ZND-conforming regime—which denotes a significant departure from “standard” detonation convergence properties—increases not solely with a reduction in chemical timescales, but with a reduction in the ratio of chemical to induction timescales, where the induction timescale is defined as the time taken for the CJ detonation to travel through the induction length. This ratio is provided in the last row of Table 1. Note that through activation energy modifications, the simplified cases 1 through 3 reduce both chemical timescale and this ratio. Meanwhile, ambient pressure increases in the the detailed cases reduce the chemical timescales, but not the ratio of chemical to induction timescales.
It is hypothesized that the deviation in the ZND-conforming regime observed by cases 2 and 3 in the simplified kinetics configuration is a direct consequence of (a) operator splitting errors due to the global time integration formulation, (b) chemical time integration errors that arise in highly stiff regimes, or (c) a combination of the two. In any case, since the timescale effects are not as pronounced in the detailed chemistry case, the trends in Fig. 5 highlight the importance of the ratio of the chemical timescale to the induction timescales in assessment of detonation convergence. The practical implication here is that, if the induction timescale can be maintained (i.e., if induction lengths and CJ speeds are not significantly altered), a reduction in chemical timescales due to mechanism simplifications and re-designs via combustion modeling can result in vastly improved detonation convergence properties.

### v2

@@ Abstract
Numerical simulations of detonation-containing flows have emerged as crucial tools for designing next-generation power and propulsion devices. As these tools mature, it is important for the combustion community to properly understand and isolate grid resolution effects when simulating detonations. To this end, the objective of this work is to provide a comprehensive analysis of the numerical convergence of unsteady detonation simulations, with focus on isolating the impacts of chemical timescale modifications on convergence characteristics in the context of operator splitting. With the aid of an AMReX-based adaptive mesh refinement flow solver [1]—which enables resolutions up to $\mathcal{O}(1000)$ cells-per-induction length—the convergence analysis is conducted using two kinetics configurations: (1) the simplified three-step Arrhenius-based model mechanism of Short and Quirk [2], where chemical timescales in the detonation are modified by adjusting activation energies in the initiation and branching reactions, and (2) the detailed hydrogen-air mechanism of Shepherd et al. [3], where the chemical timescales are adjusted by varying the ambient pressure. The convergence of unsteady self-sustained detonations in one-dimensional channels is then analyzed with reference to steady-state theoretical baseline solutions using these mechanisms. The goal of the analysis is to provide a detailed comparison of the effects of grid resolution on both macroscopic (peak pressures and wave speeds) and microscopic (wave structure) quantities of interest, drawing connections between the deviations from steady-state baselines and minimum chemical timescales. In particular, chemical timescale reductions were found to have minimal impact on the convergence of macroscopic properties. However, analyses of microscopic convergence trends, particularly in the reaction front location, revealed a key insight: maintaining the induction time while eliminating prohibitive chemical timescales through mechanism simplifications and combustion modeling can significantly enhance detonation convergence properties. Ultimately, this work uncovers resolution-dependent unsteady detonation convergence regimes and highlights the important role played by not only the chemical timescales, but also the ratio between the chemical timescale and induction time on the numerical convergence of the detonation wave structure.
Keywords: Detonations; Convergence; Chemical Kinetics; Chemical Timescales; Adaptive Mesh Refinement; Compressible Reacting Flow
@@ 1 Introduction
Numerical simulation of detonation waves in both canonical [4, 5, 6] and complex configurations [7, 8] is becoming increasingly important due to emerging interest in detonation-based propulsion devices [9, 10]. Such simulations are also crucial in other related applications, including condensed-phase explosives [11, 12] and hazard prediction [13, 14].
Regarding canonical simulations, significant ongoing efforts have been put toward performing detonation simulations with the aim of extracting fundamental physical insights into detonation-flow-chemistry interactions at levels beyond what is achievable with experimental diagnostics. These efforts span from one-dimensional to three-dimensional configurations. For example, one-dimensional numerical studies revolve around investigating and modeling nonlinearities (e.g., chaotic wave propagation behavior and chemistry effects) in detonations using both simplified and detailed chemical mechanisms in highly-resolved steady [15] and unsteady [16, 17, 18, 2, 19] settings.
Building on one-dimensional studies, numerical investigations in higher spatial dimensions have explored the fundamental physical properties of cellular detonation wave structures and their dynamic behavior across a wide range of fuel types and configurations. Earlier studies analyzed the unsteady behavior of detonations under operating conditions relevant to rotating [20, 21] and pulse [22] detonation engines. Since then, the rapid increase in computational capability in recent years has enabled detailed investigations of complex multi-physics interactions in canonical detonation-containing flows. Recent examples include, but are not limited to, (1) investigating the effect of liquid fuel injection on the detonation wave structure and particle distributions [23, 24], (2) capturing the effects of turbulence interactions on the detonation wave structure and propagation behavior (e.g., via imposition of fuel stratification [25, 26] or turbulence inflow generation [27]), (3) detailed investigations into the highly transient deflagration-to-detonation phenomenon [28, 29], and (4) resolved simulations of oblique detonations [30, 31].
The above canonical simulations are used to inform and supplement full-geometry simulations of detonation-based power and propulsion concepts, including pulse [32, 33], rotating [34, 35] and linear [36] detonation engines. In the space of rotating detonation engines specifically, state-of-the-art computational tools can now be used to conduct resolved, full-scale multi-phase detonation simulations of realistic geometries and fuels [8, 37]. Ultimately, in all practical detonation-based applications, detonation dynamics are complicated by counter-propagating waves [36], heterogeneous mixtures [38, 39], and surface stabilization effects [40], leading to complex detonation structures influenced by fine-scale flow-chemistry interactions that are difficult to probe experimentally due to the extreme conditions and short timescales in these systems [7].
Due to the inherent complexity and importance in achieving physically-accurate detonation simulations, significant effort has been devoted to developing robust, predictive, and scalable numerical methods for solving the governing compressible reacting Navier-Stokes equations. Recent approaches have utilized numerics rooted in finite volume methods for full-geometry [41] and canonical adaptive mesh refinement (AMR)-based simulations [42]. Additionally, higher-order extensions such as the piecewise parabolic method coupled with conservative shock tracking strategies [43], as well as discontinuous Galerkin methods [44], have been effectively used to study fundamental detonation dynamics. Alternative strategies like the ghost fluid method can also be used [45], which resolves shock-chemistry interactions through level set equations.
@@
Regardless of the approach used, proper understanding of grid convergence properties in PDE-based simulation of detonating flows is of utmost importance. Analysis of detonation convergence properties to this end has primary focused on identifying onsets of instability, with observations showing how acoustic waves generated in the induction zone at high resolutions—which would otherwise have been implicitly filtered at low resolutions—lead to detonation quenching behavior [47, 48]. Additionally, convergence analysis performed in Ref. [49] for ethylene detonations reports small variation in mean wave speeds, as well as the emergence of high-frequency and low-amplitude propagation modes in highly resolved cases. Similar behavior was observed in the context of n-Heptane detonations [50], where an initial convergence analysis step was carried out to isolate resolution effects on pulsating frequencies.
@@
As such, the goal of this work is to investigate the impact of chemical timescale modifications on the numerical convergence of unsteady detonations using simulations based on operator splitting. The objective is both to conduct a comprehensive detonation convergence analysis, and to isolate the chemical timescale as a tunable parameter from two different angles: (1) through direct modification of reaction mechanism parameters (i.e., activation energies), and (2) through modification of the simulation operating conditions (ambient pressures). The former leverages a simplified model three-step mechanism [2] and the latter leverages a detailed hydrogen mechanism [3]. To efficiently conduct numerical convergence studies, this work leverages a recently-developed AMReX-based flow solver [1], which discretizes the compressible reacting Navier-Stokes equations using a block-structured adaptive mesh refinement strategy. With this solver, simulation convergence trends in terms of quantities of interest (QoIs) relevant to detonation engine design, such as wave speeds, peak pressures, and detonation wave structures, are analyzed using highly resolved simulations of unsteady detonations. The specific contributions of this work are provided below.
  • 1.
    Detonation convergence analyses are performed using both simplified and detailed chemistry mechanisms in the AMR-based flow solver. The effects of chemical timescales on convergence are accessed through (a) activation energy modifications in the simplified mechanism, and (b) ambient pressure modifications in the detailed mechanism.
  • 2.
    Where applicable, comparisons are made to steady-state Zeldovich-von Neumann-Döring (ZND) baseline solutions, such that the effects of chemical timescale modifications on convergence trends can be monitored through deviations of QoIs from ZND solutions in highly-resolved regimes.
  • 3.
    Convergence analysis is performed for both macroscopic QoIs and the fine-scale detonation wave structure. The wave-structure analysis reveals three key convergence regimes: an under-resolved regime, a strongly-coupled regime characterized by shortened induction zones, and a ZND-conforming regime.
The paper proceeds as follows. In Sec. 2, the numerical approach is outlined, and details are provided for the chemical mechanisms and the processes by which the chemical timescales are modified. Results are provided in Sec. 3, with Sec. 3.1 outlining the steady-state solution baselines, Sec. 3.2 detailing the convergence trends of macroscopic detonation QoIs (e.g., peak pressures and wave speeds), and Sec. 3.3 providing convergence trends in terms of the fine-scale detonation wave structure. Concluding remarks are provided in Sec. 4.
@@ 2 Numerical Approach
The governing equations are the compressible reacting Navier-Stokes equations:
$\frac{\partial\rho}{\partial t}+\frac{\partial\rho u_{i}}{\partial x_{i}}=0,$ (1)
$\frac{\partial\rho u_{j}}{\partial t}+\frac{\partial\rho u_{i}u_{j}}{\partial x_{i}}=-\frac{\partial p}{\partial x_{j}}+\frac{\partial\tau_{ij}}{\partial x_{i}},$ (2)
$\frac{\partial\rho E}{\partial t}+\frac{\partial\rho u_{i}H}{\partial x_{i}}=\frac{\partial}{\partial x_{i}}\left(\alpha\frac{\partial T}{\partial x_{i}}\right)+\frac{\partial\tau_{ij}u_{i}}{\partial x_{j}}+\sum_{k}\left(h_{k}\frac{\partial}{\partial x_{i}}\left(\rho D\frac{\partial Y_{k}}{\partial x_{i}}\right)\right),$ (3)
$\frac{\partial\rho Y_{k}}{\partial t}+\frac{\partial\rho u_{i}Y_{k}}{\partial x_{i}}=\frac{\partial}{\partial x_{i}}\left(\rho D\frac{\partial Y_{k}}{\partial x_{i}}\right)+\dot{\Omega}_{k},$ (4)
where
$\tau_{ij}=-\frac{2}{3}\mu\frac{\partial u_{k}}{\partial x_{k}}\delta_{ij}+\mu\left(\frac{\partial u_{j}}{\partial x_{i}}+\frac{\partial u_{i}}{\partial x_{j}}\right).$ (5)
Equations 1-4 respectively correspond to mass, momentum, energy, and species conservation. Here, $\rho$ is the fluid density, $t$ is time, $x_{i}$ is the spatial position, $u_{i}$ is the fluid velocity, $p$ is the fluid pressure, $\tau_{ij}$ is the viscous stress tensor, $\mu$ is the dynamic viscosity, $E$ is the total chemical energy, $H$ is the total mixture enthalpy, $h_{k}$ is the species enthalpy (including sensible and chemical components), $\alpha$ is the thermal conductivity, $Y_{k}$ is the mass fraction of the $k$-th species, $D$ is the local diffusivity (assumed equal for all species), and $\dot{\Omega}_{k}$ is the chemical source term for species $k$. In this work, viscosity and diffusion effects are neglected in the governing equations for both the simplified and detailed chemical mechanisms. An ideal gas equation of state is nominally used to close the system. In the simplified mechanism detailed in Sec. 2.1, the gas is further assumed to be calorically perfect.
[…]
The equations are solved using an in-house solver based on the AMReX framework [1]. The solver is a block-structured adaptive mesh refinement (AMR) extension of the extensively verified UMReactingFlow [41]. On each mesh refinement level, chemical time integration is handled using an adaptive explicit method, and the advection-diffusion advance is treated using a stability-preserving second-order Runge-Kutta method combined with a slope-limited approximate Riemann solver. This work specifically leverages a van Leer limiter with the Harten–Lax–van Leer contact (HLLC) approximate Riemann solver [52]. Grid refinement is handled by AMReX [53], which provides a backend for the grid communication steps used in the block-structured AMR formulation. The reader is directed to Ref. [1] for additional detail on the numerics of each component and the solution procedure.
The AMR tagging criterion is based on pressure and density gradients with additional buffer cell padding, and was determined to properly isolate the detonation wave structures for subsequent mesh refinement. Grid refinement factors were set to 2, and all convergence resolutions reported in Sec. 3 are extracted from the finest (highest) AMR level, with the highest level count being 14. Separate verification studies were conducted to ensure that AMR-induced interpolation errors did not pollute detonation profiles for the range of levels considered. Examples of unsteady detonation profiles showcasing AMR level coverage in a 12-level case for both simplified (Sec. 2.1) and detailed (Sec. 2.2) chemistry configurations is shown in Fig. 1. The figure shows that the refinement criteria used in the AMR simulations successfully resolve the detonation profile at the desired level specification.
@@ 2.1 Simplified Chemistry
There are four key parameters: the initiation and branching activation energies ($E_{I}$ and $E_{B}$), and the corresponding crossover temperatures ($T_{I}$ and $T_{B}$). This work leverages the activation energies to control the sensitivity of the detonating flow field to chemical reactions, providing a direct pathway to parameterize chemical timescales within the detonation wave structure, as showcased in Sec. 3.1. In particular, three different activation energy settings (Cases 1-3 in Table 2) were prescribed to explore a wide range of chemical timescale influence in the simplified configurations. Besides the activation energies, all other parameters match those used for the lowest $T_{B}$ case in Ref. [54]. To admit a theoretically valid detonation solution with this mechanism, the gas is assumed to be calorically perfect with constant specific heat ratio $\gamma$, leading to a simplified definition of chemical energy. This differs from the detailed case, where variable $\gamma$ is used. All length and time values in the simplified configurations are reported in non-dimensional quantities, where non-dimensionalization is carried out such that the termination rate $k_{C}$ is unity. The reader is directed to Ref. [54] for additional details.
@@ 2.2 Detailed Chemistry
Detailed chemical kinetics are modeled using the 14 species, 42 reaction H2-air mechanism of Shepherd et al. [3], which is based on the mechanism of Mével et al. [55]  [^1] The mechanism is openly available at the following link: https://shepherd.caltech.edu/EDL/PublicResources/sdt/SDToolbox/cti/Mevel2017.cti. The one-dimensional domain was 0.8 m long with a uniform base grid of 1600 cells, yielding a base resolution of 500 $\mu$m. The minimum grid size in the finest simulation is 30.5 nm. All cases used the reactant molar composition 2H2:O2:7Ar at an ambient temperature of 300 K. Significant argon dilution was added to ensure stability of the detonation waves [56]. In order to explore different chemical timescale regimes, ambient pressures of 0.2, 0.5, and 1 atm were investigated (as described in Table 2). For each case, the ZND solution was computed using the Shock and Detonation Toolbox [57]. To initialize unsteady simulations, this solution was then mapped onto the computational grid at the initial time, with the wave front positioned 0.2 m from the left boundary (i.e., the driver gas was the ZND wave in the detailed cases, as these were found to more quickly lead to self-sustained detonations).
Symbol Description
$E_{I},E_{B}$ Initiation and branching activation energies used in the simplified mechanism.
$T_{I},T_{B}$ Initiation and branching crossover temperatures using in the simplified mechanism.
$c_{0}$ Ambient speed-of-sound.
$P_{0}$ Ambient pressure.
$\tau_{\text{chem}}$ Minimum chemical timescale observed in the detonation wave, obtained from the largest eigenvalue of the chemical source term Jacobian.
$\tau_{\text{ind}}$ Induction time.
$L_{\text{ind}}$ Induction zone length.
$L_{\text{react}}$ Reaction zone length.
$W$ Detonation wave speed observed from numerical simulations.
$W_{\text{CJ}}$ Wave speed of Chapman-Jouguet detonation.
$P_{\text{max}}$ Maximum pressure observed on the detonation wave.
$P_{\text{VN}}$ Von-Neumann pressure observed in the theoretical ZND solution.
$x_{s}$ Spatial position relative to leading shockwave.
Table 1: Reference nomenclature table for relevant symbols and variables used throughout this paper.
@@ 3 Results
$\tau_{\text{chem}}/\tau_{\text{ind}}$ 2.333 ×$\times 10^{-2}$ 3.645 ×$\times 10^{-4}$ 1.033 ×$\times 10^{-5}$ 3.680 ×$\times 10^{-4}$ 3.985 ×$\times 10^{-4}$ 4.242 ×$\times 10^{-4}$
Table 2: Detonation case configurations. First row: activation energies in simplified mechanism normalized by squared speed-of-sound. Second row: ambient pressures. Third row: minimum chemical timescale observed in steady-state solution. Fourth row: Ratio of minimum chemical timescale and induction time in steady-state solution. Induction time defined as $\tau_{\text{ind}}=\int_{0}^{L_{\text{ind}}}\frac{1}{U(x_{s})}\text{d}x_{s}$, where $L_{\text{ind}}$ is induction length, $x_{s}$ distance behind the leading shock, and $U(x_{s})$ is the local fluid velocity in the shock reference frame. ZND profiles shown in Fig. 2.
@@ 3.3 Convergence of Wave Structure
This finding illuminates how deviation from the steady-state solution in the ZND-conforming regime—which denotes a significant departure from “standard” detonation convergence properties—increases not solely with a reduction in chemical timescales, but with a reduction in the ratio of chemical timescales to induction times, where the induction time is defined as the integrated inverse local fluid velocity in the induction zone (see caption in Table 2). This ratio is provided in the last row of Table 2. Note that through activation energy modifications, the simplified cases 1 through 3 reduce both chemical timescale and this ratio. Meanwhile, ambient pressure increases in the the detailed cases reduce the chemical timescales, but not the ratio of chemical timescale to induction time.
It is hypothesized that the deviation in the ZND-conforming regime observed by cases 2 and 3 in the simplified kinetics configuration is a direct consequence of (a) operator splitting errors due to the global time integration formulation, (b) chemical time integration errors that arise in highly stiff regimes, or (c) a combination of the two. In any case, since the timescale effects are not as pronounced in the detailed chemistry case, the trends in Fig. 5 highlight the importance of the ratio of the chemical timescale to the induction times in assessment of detonation convergence. The practical implication here is that, if the induction time can be maintained (i.e., if induction lengths and CJ speeds are not significantly altered), a reduction in chemical timescales due to mechanism simplifications and re-designs via combustion modeling can result in vastly improved detonation convergence properties.